OpenStudy (fibonaccichick666):

One card is selected from a standard deck of 52 cards. What is the probability that the card is either a diamond or a face card?

Nnesha (nnesha):

there are 13 face cards right :P o,O hmm

OpenStudy (fibonaccichick666):

Looking for others' interpretations here. I thought it should be \(\frac{13}{52}+\frac{9}{52}=24/52=12/26=6/13\)

OpenStudy (fibonaccichick666):

I'm reading the problem wrong somehow(or book has a type-o)

OpenStudy (fibonaccichick666):

Yea, 13 face cards nosh

Nnesha (nnesha):

and 13 diamonds lol ??? no idea abt cards :P

OpenStudy (amistre64):

some rules define aces as faces

OpenStudy (fibonaccichick666):

They say the answer is 13/52

OpenStudy (fibonaccichick666):

Made that mistake ( I defined an ace as a face first time around)

OpenStudy (fibonaccichick666):

but this is my second go

OpenStudy (fibonaccichick666):

Shouldn't the probability be greater than just diamonds if we have an or and not an XOR?

OpenStudy (amistre64):

if we remove all the face cards, and the diamonds, that is our ... outcome space?

OpenStudy (amistre64):

10+12 = 22 cards that are favored, out of 52 total

OpenStudy (amistre64):

1thru9 jkq 9+12 = 18 ... out of 52 :)

OpenStudy (fibonaccichick666):

wait, only KQJ are face cards.... I think

OpenStudy (amistre64):

XOR doesnt seem to translate well into english prose

OpenStudy (amistre64):

its 11pm .. and i forgot how to add ... im going to bed :)

OpenStudy (fibonaccichick666):

lol, I am a bit confused by your argument. Let me attempt to understand

OpenStudy (fibonaccichick666):

we should have 30 cards not an option

ganeshie8 (ganeshie8):



OpenStudy (fibonaccichick666):

3 suits, A-10 each

OpenStudy (fibonaccichick666):

I added wrong above My conclusion should be 11/26

Nnesha (nnesha):

told you then i deleted the comment i thought this is impossible how fib can got it wrong lol

OpenStudy (fibonaccichick666):

lol I can't add today, clearly

OpenStudy (fibonaccichick666):

but still they are claiming 13/52 is the answer

OpenStudy (fibonaccichick666):

@ganeshie8 Would you say 13/52 is incorrect?

OpenStudy (fibonaccichick666):

I mean I think you proved it, but just want confirmation

ganeshie8 (ganeshie8):

A : diamonds B : face cards n(A) = 13 n(B) = 12 n(A \(\cap\) B) = 3 outcomes in favor = n(A \(\cup \) B) = n(A) + n(B) - n(A \(\cap\) B) = 13 + 12 - 3 = 22 total outcomes = 52 so p(A \(\cup\) B) = 22/52

OpenStudy (fibonaccichick666):

here is their argument: "The prob. that a card is either a diamond or a face card is determined as follows: Let D be the event the card is a diamond. Let F be the event the card is a face card. Then, Pr[D}=13/52 and Pr[F]=12/52. We want to compute the probability of the event DUF Pr[DUF]=Pr[D]+Pr[F]-Pr[D(intersect)F]=13/52+12/52-3/13=13/52

OpenStudy (fibonaccichick666):

Ok, so book is def. wrong and I can't add... Great. Thanks guys!